# Information system for analysis of nanostructure morphology

### v1

@@ Abstract
The primary focus of this work is the automating of morphology analysis for colloidal structures, achieved through the development of the ISANM information system. The ISANM system integrates various analysis methods and is designed prioritize usability, catering primarily to users rather than developers. This paper outlines the scope of the subject area, reviews selected methods for morphology analysis, and presents the scripts developed for implementing these methods in Python. The system’s core is implemented in C#, utilizing the .NET Framework and the MS SQL Server, and is developed within the Microsoft Visual Studio 2019 environment on Windows 10 using the client-server architecture. The system has been deployed on a VDS server using the Ubuntu 20.04 distribution. MS SQL Server is used as the database management system. The system can be accessed at https://isanm.space. An ORCID account is required for authorization. Geometric analysis of colloidal structures is significant in applications such as photonic crystals for optoelectronics and microelectronics, functional coatings in materials science, and biosensors for medical and environmental applications.
@@
information system , micro- and nanoparticles , morphology of structures , analysis methods , database
journal: Information systems
@@ Introduction
Some existing software solutions support various methods for analyzing colloidal morphology. For instance, ImageJ and Image-Pro Plus are used for image analysis and processing, while MATLAB is utilized for data processing. Another example, MATBOX is an open-source toolbox for the commercial software package MATLAB [7]. MATBOX is the application for microstructure numerical analysis, including microstructure numerical generation, image filtering and microstructure segmentation, microstructure characterization, three-dimensional visualization, microstructure parameters correlation, and microstructure meshing. The JavaPlex library implements persistent homology and related techniques from the field of computational and applied topology. Generally, such software is more oriented towards experienced users or programmers. One of the advantages of ISANM is the integration of various analysis methods within a single, accessible service, that does not require local installation. Data storage and processing occur on the server, and only basic computer skills are necessary, making it accessible to engineers and scientists working with colloidal structures and their applications. Geometric analysis of colloidal assemblies is important in areas such as photonic crystals for opto- and microelectronics, functional coatings in materials science, and the creation of biosensors for medical and environmental applications.
@@ 1.3 Radial uniformity
Radial uniformity describes the property of a material or system that tends to exhibit uniform characteristics in the radial direction relative to a central point, axis, or center of mass. This concept can be applied in various fields, such as physics, engineering, or geometry. In the context of colloidal structures, radial uniformity may refer to the uniform distribution of colloidal particles relative to the center of the colloidal system. Colloidal systems include particles ranging in size from a few nanometers to micrometers, dispersed within a medium such as water. Assessment of the radial uniformity of colloidal particles may include measurement and analysis of particle concentrations in various radial zones or distances from the center of a colloidal system.
@@ 2.4 User authorization in the system
When the system is launched, users are immediately directed to the authentication page, where they need to complete the authentication process to access the full range of functionalities. The available features will vary according to the authenticated user’s role. Currently, users can authenticate using their Open Researcher and Contributor ID (ORCID), allowing users with an existing ORCID account to log into the system seamlessly. Authentication is performed via the ORCID API (see https://orcid.org/), which provides secure access to user data. This method not only streamlines the registration and login process but also enables users to leverage their existing accounts. As a result, this approach safeguards the system from unverified users while also simplifying and automating the authentication process. The full range of system functionalities available to users is depicted in the Use Case diagram shown in Fig. 18.
@@ 3 Impact
The proposed system integrates various methods for analyzing colloidal structures in one platform. This online service is beneficial for engineers and scientists working with colloidal assemblies and their applications. Using the ISANM information system for data processing allows users to automate workflows without the need for programming skills.
@@ CONCLUSION
The developed Information system for analysis of nanostructure morphology, ISANM, provides experts in the relevant field with useful tools for their work, requiring no programming skills. This system functions as a client-server application, ensuring ease of use, accessibility, and the delegation of computational tasks from the client to the server. Users can upload and store their files in the system, categorize them into thematic folders, apply various processing methods, save the results, and download output files containing the processed data. In the future, the developed tool will be useful for scientists and engineers working in the fields of colloidal chemistry, materials science and nanotechnology, which will affect the increase in labor productivity.

### v2

@@ Abstract
The work is devoted to the development of an information system ISANM that will be useful for teaching students and applicable in the work of engineers and researchers for automating the analysis of colloidal structure morphology. The ISANM system integrates various analysis methods and is designed prioritize usability, catering primarily to users rather than software developers. This paper outlines the scope of the subject area, reviews selected methods for morphology analysis, and presents the scripts developed for implementing these methods in Python. In the future, new methods and tools will be gradually added to the system for chemical engineers, physicists, and materials scientists. We hope that the system implementation methodology described here will be useful in the implementation of other projects related to training chemical engineers and beyond. The system core is implemented in C#, utilizing the .NET Framework and the MS SQL Server, and is developed within the Microsoft Visual Studio 2019 environment on Windows 10 using the client-server architecture. The system has been deployed on a VDS server using the Ubuntu 20.04 distribution. MS SQL Server is used as the database management system. The system can be accessed at https://isanm.space. Geometric analysis of colloidal structures is significant in applications such as photonic crystals for optoelectronics and microelectronics, functional coatings in materials science, and biosensors for medical and environmental applications. This demonstrates the importance of digitalization in this area to improve the quality of student education.
@@
virtual laboratories , information systems , micro- and nanoparticles , morphology of structures , analysis methods , databases , simulation tools in education
journal:
@@ Introduction
Digitalization is one of the most important issues in chemical engineering education [1, 2]. An example of a 5-year curriculum is given in one pioneering work [3]. It includes many training modules, including “Advances in colloids and interfaces”. The development of digital tools in this field is crucial for the future development of educational and research methods.
@@
Some existing software solutions support various methods for analyzing colloidal morphology. For instance, ImageJ and Image-Pro Plus are used for image analysis and processing, while MATLAB is utilized for data processing. Another example, MATBOX is an open-source toolbox for the commercial software package MATLAB [10]. MATBOX is the application for microstructure numerical analysis, including microstructure numerical generation, image filtering and microstructure segmentation, microstructure characterization, three-dimensional visualization, microstructure parameters correlation, and microstructure meshing. The JavaPlex library implements persistent homology and related techniques from the field of computational and applied topology. Generally, such software is more oriented towards experienced users or programmers. One of the advantages of ISANM is the integration of various analysis methods within a single, accessible service, that does not require local installation. Data storage and processing occur on the server, and only basic computer skills are necessary, making it accessible to engineers and scientists working with colloidal structures and their applications. In addition, the ISANM system does not compete with the mentioned programs, such as ImageJ. Rather, they complement each other and can be used together. In the future, we will think about integration with different tools. Geometric analysis of colloidal assemblies is important in areas such as photonic crystals for opto- and microelectronics, functional coatings in materials science, and the creation of biosensors for medical and environmental applications. This demonstrates the importance of digitalization in this area to improve the quality of student education.
The purpose of this work is to demonstrate the possibility of creating new digital tools for educational and research activities using ISANM development as an example. Here, we share our experience in this area, describe the methodology for implementing such a project, consider the theoretical foundations for analyzing the morphology of colloidal structures, and give instructions on how to use the implemented functionality in practice, which may be useful for educators, students, and researchers.
@@ 1.3 Radial uniformity
Radial uniformity describes the property of a material or system that tends to exhibit uniform characteristics in the radial direction relative to a central point, axis, or center of mass. This concept can be applied in various fields, such as physics, engineering, or geometry. In the context of colloidal structures, radial uniformity may refer to the uniform distribution of colloidal particles relative to the center of the colloidal system. Colloidal systems include particles ranging in size from a few nanometers to micrometers, dispersed within a medium such as water or another liquid. Assessment of the radial uniformity of colloidal particles may include measurement and analysis of particle concentrations in various radial zones or distances from the center of a colloidal system.
@@ 2.4 User authorization in the system
When the system is launched, users are immediately directed to the authentication page, where they need to complete the authentication process to access the full range of functionalities. The available features will vary according to the authenticated user’s role. Currently, users can authenticate using their Open Researcher and Contributor ID (ORCID), allowing users with an existing ORCID account to log into the system seamlessly. Authentication is performed via the ORCID API (see https://orcid.org/), which provides secure access to user data. This method not only streamlines the registration and login process but also enables users to leverage their existing accounts. As a result, this approach safeguards the system from unverified users while also simplifying and automating the authentication process. The full range of system functionalities available to users is depicted in the Use Case diagram shown in Fig. 18. If desired, after logging in for the first time, one can confirm own email address and create a username/ password to log into their personal account.
@@ 3 Impact
The proposed system integrates various methods for analyzing colloidal structures in one platform. This online service is beneficial for engineers and scientists working with colloidal assemblies and their applications. In addition, this system is useful for teachers and students, as it can be integrated into the learning process. Using the ISANM information system for data processing allows users to automate workflows without the need for programming skills. This will speed up research and improve the quality of education for chemical engineers and beyond.
@@
4 Further work
In the future, it is planned to increase the number of analysis methods in the system, including those related to image processing and based on machine learning and neural networks. The reference materials can be made even more visual and user-friendly, including new methods. It is also planned to introduce a number of mathematical models describing the self-assembly of colloidal structures. Another important area is the integration of work with other tools, such as ImageJ (data export and import, similar data formats, and so on). It may be useful to be able to share the user’s folder via a link or another method. This will allow students to carry out joint projects. They will have convenient opportunities to demonstrate their results to teachers. Teachers will be able to share their materials with students. Researchers will organize collaboration in their work. Feedback will definitely be implemented, which will allow us to make timely edits to the system and add new functionality based on user wishes.
@@ Conclusion
The developed Information system for analysis of nanostructure morphology, ISANM, provides experts in the relevant field with useful tools for their work, requiring no programming skills. This system functions as a client-server application, ensuring ease of use, accessibility, and the delegation of computational tasks from the client to the server. Users can upload and store their files in the system, categorize them into thematic folders, apply various processing methods, save the results, and download output files containing the processed data. In the future, the developed tool will be useful for scientists and engineers working in the fields of colloidal chemistry, materials science and nanotechnology, which will affect the increase in labor productivity. In our opinion, introducing this system into the educational process would improve the quality of training for future specialists.